A population of lizards is growing at a rate of 4.6% / year. Initially, there are 1750 lizards. How long will it take for the population to double? (round to two decimal places)
A.14.72 years
B.16.21 years
C.15.41 years
D.Solution not listed

Answer by Boreal(15235)

Rule of 70 (70/4.6=15.22) suggests C
3500=1750(1.046)^n
2=1.046^n
ln2=n ln 1.046
0.693/0.0445=n. The 0.693 is the decimal of 69.3%, and that is why the Rule of 70 exists for doubling. Some use 72, since so many more numbers divide evenly into it.
n=15.41 years.
C